Sum of infinite series  1/9+1/18+1/30+1/45+1/63...
How do I solve the question
Kindly give a detailed solution

Grade:12

1 Answers

vivek
7 years ago
9 18 30 45 t(n)=an^2+bn+c 9 12 15 9=a+b+c ; 18=4a+2b+c; 3 3 30=9a+3b+ca=3/2,b=9/2,c=3 for ques t(n)=2/3(1/(n+1)(n+2)sum(1 - infinity)= 2/3 (1/n+1 - 1/n+2 )=2/3 (1/2) =1/3